# 2D Gaussian Splatting for Geometrically Accurate Radiance Fields

### v1

@@ Abstract.
3D Gaussian Splatting (3DGS) has recently revolutionized radiance field reconstruction, achieving high quality novel view synthesis and fast rendering speed. However, 3DGS fails to accurately represent surfaces due to the multi-view inconsistent nature of 3D Gaussians. We present 2D Gaussian Splatting (2DGS), a novel approach to model and reconstruct geometrically accurate radiance fields from multi-view images. Our key idea is to collapse the 3D volume into a set of 2D oriented planar Gaussian disks. Unlike 3D Gaussians, 2D Gaussians provide view-consistent geometry while modeling surfaces intrinsically. To accurately recover thin surfaces and achieve stable optimization, we introduce a perspective-accurate 2D splatting process utilizing ray-splat intersection and rasterization. Additionally, we incorporate depth distortion and normal consistency terms to further enhance the quality of the reconstructions. We demonstrate that our differentiable renderer allows for noise-free and detailed geometry reconstruction while maintaining competitive appearance quality, fast training speed, and real-time rendering. Our code will be made publicly available.
@@
journal: TOGccs: Computing methodologies Reconstructionccs: Computing methodologies Renderingccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
On the other hand, earlier works (Pfister et al., 2000; Zwicker et al., 2001a) have shown surfels (surface elements) to be an effective representation for complex geometry. Surfels approximate the object surface locally with shape and shade attributes, and can be derived from known geometry. They are widely used in SLAM (Whelan et al., 2016) and other robotics tasks (Schöps et al., 2019) as an efficient geometry representation. Subsequent advancements (Yifan et al., 2019) have incorporated surfels into a differentiable framework. However, these methods typically require ground truth (GT) geometry, depth sensor data, or operate under constrained scenarios with known lighting.
Inspired by these works, we propose 2D Gaussian Splatting for 3D scene reconstruction and novel view synthesis that combines the benefits of both worlds, while overcoming their limitations. Unlike 3DGS, our approach represents a 3D scene with 2D Gaussian primitives, each defining an oriented elliptical disk. The significant advantage of 2D Gaussian over its 3D counterpart lies in the accurate geometry representation during rendering. Specifically, 3DGS evaluates a Gaussian’s value at the intersection between a pixel ray and a 3D Gaussian (Keselman and Hebert, 2022, 2023), which leads to inconsistency depth when rendered from different viewpoints. In contrast, our method utilizes explicit ray-splat intersection, resulting in a perspective accuracy splatting, as illustrated in Figure 2, which in turn significantly improves reconstruction quality. Furthermore, the inherent surface normals in 2D Gaussian primitives enable direct surface regularization through normal constraints. In contrast with surfels-based models (Pfister et al., 2000; Zwicker et al., 2001a; Yifan et al., 2019), our 2D Gaussians can be recovered from unknown geometry with gradient-based optimization.
@@
  • We present a highly efficient differentiable 2D Gaussian renderer, enabling perspective-accurate splatting by leveraging 2D surface modeling, ray-splat intersection, and volumetric integration.
@@ 2.1. Novel view synthesis
3D Reconstruction from multi-view images has been a long-standing goal in computer vision. Multi-view stereo based methods (Schönberger et al., 2016; Yao et al., 2018; Yu and Gao, 2020) rely on a modular pipeline that involves feature matching, depth prediction, and fusion. In contrast, recent neural approaches (Niemeyer et al., 2020; Yariv et al., 2020) represent surface implicilty via an MLP (Park et al., 2019; Mescheder et al., 2019) , extracting surfaces post-training via the Marching Cube algorithm. Further advancements (Oechsle et al., 2021; Wang et al., 2021; Yariv et al., 2021) integrated implicit surfaces with volume rendering, achieving detailed surface reconstructions from RGB images. These methods have been extended to large-scale reconstructions via additional regularization (Yu et al., 2022b; Li et al., 2023; Yu et al., 2022a), and efficient reconstruction for objects (Wang et al., 2023). Despite these impressive developments, efficient large-scale scene reconstruction remains a challenge. For instance, Neuralangelo (Li et al., 2023) requires 128 GPU hours for reconstructing a single scene from the Tanks and Temples Dataset (Knapitsch et al., 2017). In this work, we introduce 2D Gaussian splatting, a method that significantly accelerates the reconstruction process. It achieves similar or slightly better results compared to previous implicit neural surface representations, while being an order of magnitude faster.
@@ 2.4. Concurrent work
Challenges in Surface Reconstruction: Reconstructing surfaces using 3D Gaussian modeling and splatting faces several challenges. First, the volumetric radiance representation of 3D Gaussians conflicts with the thin nature of surfaces. Second, 3DGS does not natively model surface normals, essential for high-quality surface reconstruction. Third, the rasterization process in 3DGS lacks multi-view consistency, leading to varied 2D intersection planes for different viewpoints (Keselman and Hebert, 2023), as illustrated in Figure 2 (a). Additionally, using an affine matrix for transforming a 3D Gaussian into ray space only yields accurate projections near the center, compromising on perspective accuracy around surrounding regions (Zwicker et al., 2004). Therefore, it often results in noisy reconstructions, as shown in Figure 5.
@@ 4.1. Modeling
Unlike 3DGS (Kerbl et al., 2023), which models the entire angular radiance in a blob, we simplify the 3-dimensional modeling by adopting “flat” 2D Gaussians embedded in 3D space. With 2D Gaussian modeling, the primitive distributes densities within a planar disk, defining the normal as the direction of steepest change of density. This feature enables better alignment with thin surfaces. While previous methods (Kopanas et al., 2021; Yifan et al., 2019) also utilize 2D Gaussians for geometry reconstruction, they require a dense point cloud or ground-truth normals as input. By contrast, we simultaneously reconstruct the appearance and geometry given only a sparse calibration point cloud and photometric supervision.
@@
(4) $P(u,v)=\mathbf{p}_{k}+s_{u}\mathbf{t}_{u}u+s_{v}\mathbf{t}_{v}v=\mathbf{H}(u,v,1,1)^{\top}$
@@ 4.2. Splatting
One common strategy for rendering 2D Gaussians is to project the 2D Gaussian primitives onto the image space using the affine approximation for the perspective projection (Zwicker et al., 2001a). However, as noted in  (Zwicker et al., 2004), this projection is only accurate at the center of the Gaussian and has increasing approximation error with increased distance to the center. To address this issue, Zwicker et al. proposed a formulation based on homogeneous coordinates. Specifically, projecting the 2D splat onto an image plane can be described by a general 2D-to-2D mapping in homogeneous coordinates. Let $\mathbf{W}\in{4\times 4}$ be the transformation matrix from world space to screen space. The screen space points are hence obtained by
(7) $\mathbf{x}=(xz,yz,z,z)^{\top}=\mathbf{W}P(u,v)=\mathbf{W}\mathbf{H}(u,v,1,1)^{\top}$
@@
Ray-splat Intersection: We efficiently locate the ray-splat intersections by finding the intersection of three non-parallel planes, a method initially designed for specialized hardware (Weyrich et al., 2007). Given an image coordinate $\mathbf{x}=(x,y)$, we parameterize the ray of a pixel as the intersection of two orthogonal planes: the x-plane and the y-plane. Specifically, the x-plane is defined by a normal vector $(-1,0,0)$ and an offset $x$. Therefore, the x-plane can be represented as a 4D homogeneous plane $\mathbf{h}_{x}=(-1,0,0,x)$. Similarly, the y-plane is $\mathbf{h}_{y}=(0,-1,0,y)$. Thus, the ray $\mathbf{x}=(x,y)$ is determined by the intersection of the $x$-plane and the $y$-planes.
Next, we transform both planes into the local coordinates of the 2D Gaussian primitives, the $uv$-coordinate system. Note that transforming points on a plane using a transformation matrix $\mathbf{M}$ is equivalent to transforming the homogeneous plane parameters using the inverse transpose $\mathbf{M}^{-\top}$ (Vince, 2008). Therefore, applying $\mathbf{M}=(\mathbf{W}\mathbf{H})^{-1}$ is equivalent to $(\mathbf{W}\mathbf{H})^{\top}$, eliminating explicit matrix inversion and yielding:
(8) $\mathbf{h}_{u}=(\mathbf{W}\mathbf{H})^{\top}{\mathbf{h}_{x}}\quad\mathbf{h}_{v}=(\mathbf{W}\mathbf{H})^{\top}{\mathbf{h}_{y}}$
@@
(9) $\mathbf{h}_{u}\cdot(u,v,1,1)^{\top}=\mathbf{h}_{v}\cdot(u,v,1,1)^{\top}=0$
@@
where $\mathbf{h}_{u}^{i},\mathbf{h}_{v}^{i}$ are the $i$-th parameters of the 4D homogeneous plane parameters. We obtain the depth $z$ of the intersected points via Eq. 7 and evaluate the Gaussian value with Eq 6.
Degenerate Solutions: When a 2D Gaussian is observed from a slanted viewpoint, it degenerates to a line in screen space. Therefore, it might be missed during rasterization. To deal with these cases and stabilize optimization, we employ the object-space low-pass filter introduced in (Botsch et al., 2005):
@@
where $\mathbf{u}(\mathbf{x})$ is given by (10) and $\mathbf{c}$ is the projection of center $\mathbf{p}_{k}$. Intuitively, $\hat{\mathcal{G}}(\mathbf{x})$ is lower-bounded by a fixed screen-space Gaussian low-pass filter with center $\mathbf{c}_{k}$ and radius $\sigma$. In our experiments, we set $\sigma=\sqrt{2}/2$ to ensure sufficient pixels are used during rendering.
Rasterization: We follow a similar rasterization process as in 3DGS (Kerbl et al., 2023). First, a screen space bounding box is computed for each Gaussian primitive. Then, 2D Gaussians are sorted based on the depth of their center and organized into tiles based on their bounding boxes. Finally, volumetric alpha blending is used to integrate alpha-weighted appearance from front to back:
@@
Figure 4. Visual comparisons (test-set view) between our method, 3DGS, and SuGaR using scenes from an unbounded real-world dataset - kitchen, bicycle, stump, and treehill. The first column displays the ground truth novel view, while the second and the third columns showcase our rendering results, demonstrating high-fidelity novel view and accurate geometry. Columns four and five present surface reconstruction results from our method and 3DGS, utilizing TSDF fusion with a depth range of 6 meters. Background regions (outside this range) are colored light gray. The sixth column displays results from SuGaR. In contrast to both baselines, our method excels in capturing sharp edges and intricate geometry.
@@ 5. Training
Depth Distortion: Different from NeRF, 3DGS’s volume rendering doesn’t consider the distance between intersected Gaussian primitives. Therefore, spread out Gaussians might result in a similar color and depth rendering. This is different from surface rendering, where rays intersect the first visible surface exactly once. To mitigate this issue, we take inspiration from Mip-NeRF360 (Barron et al., 2022a) and propose a depth distortion loss to concentrate the weight distribution along the rays by minimizing the distance between the ray-splat intersections:
@@
2DGS-30k (Ours) 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80 18.8 m
Normal Consistency: As our representation is based on 2D Gaussian surface elements, we must ensure that all 2D splats are locally aligned with the actual surfaces. In the context of volume rendering where multiple semi-transparent surfels may exist along the ray, we consider the actual surface at the median point of intersection, where the accumulated opacity reaches 0.5. We then align the splats’ normal with the gradients of the depth maps as follows:
(14) $\mathcal{L}_{n}=\sum_{i}\omega_{i}(1-\mathbf{n}_{i}^{\top}\mathbf{N})$
where $i$ indexes over intersected splats along the ray, $\omega$ denotes the blending weight of the intersection point, $\mathbf{n}_{i}$ represents the normal of the splat that is oriented towards the camera, and $\mathbf{N}$ is the normal estimated by the nearby depth point $\mathbf{p}$. Specifically, $\mathbf{N}$ is computed with finite differences as follows:
(15) $\mathbf{N}(x,y)=\frac{\nabla_{x}\mathbf{p}\times\nabla_{y}\mathbf{p}}{|\nabla_{x}\mathbf{p}\times\nabla_{y}\mathbf{p}|}$
@@
Final Loss: Finally, we optimize our model from an initial sparse point cloud using a set of posed images. We minimize the following loss function:
@@
where $\mathcal{L}c$ is an RGB reconstruction loss combining $\mathcal{L}_{1}$ with the D-SSIM term from (Kerbl et al., 2023), while $\mathcal{L}_{d}$ and $\mathcal{L}_{n}$ are regularization terms. We set $\alpha=1000$ for bounded scenes, $\alpha=100$ for unbounded scenes and $\beta=0.05$ for all scenes.

### v2

@@ Abstract.
3D Gaussian Splatting (3DGS) has recently revolutionized radiance field reconstruction, achieving high quality novel view synthesis and fast rendering speed. However, 3DGS fails to accurately represent surfaces due to the multi-view inconsistent nature of 3D Gaussians. We present 2D Gaussian Splatting (2DGS), a novel approach to model and reconstruct geometrically accurate radiance fields from multi-view images. Our key idea is to collapse the 3D volume into a set of 2D oriented planar Gaussian disks. Unlike 3D Gaussians, 2D Gaussians provide view-consistent geometry while modeling surfaces intrinsically. To accurately recover thin surfaces and achieve stable optimization, we introduce a perspective-accurate 2D splatting process utilizing ray-splat intersection and rasterization. Additionally, we incorporate depth distortion and normal consistency terms to further enhance the quality of the reconstructions. We demonstrate that our differentiable renderer allows for noise-free and detailed geometry reconstruction while maintaining competitive appearance quality, fast training speed, and real-time rendering.
@@
copyright: acmlicensedjournalyear: 2024copyright: rightsretainedconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24; July 27-August 1, 2024; Denver, CO, USAbooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24 (SIGGRAPH Conference Papers ’24), July 27-August 1, 2024, Denver, CO, USAdoi: 10.1145/3641519.3657428isbn: 979-8-4007-0525-0/24/07ccs: Computing methodologies Reconstructionccs: Computing methodologies Renderingccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
On the other hand, earlier works (Pfister et al., 2000; Zwicker et al., 2001a, b) have shown surfels (surface elements) to be an effective representation of complex geometry. Surfels approximate the object surface locally with shape and shade attributes and can be derived from known geometry. They are widely used in SLAM (Whelan et al., 2016) and other robotics tasks (Schöps et al., 2019) as an efficient geometry representation. Subsequent advancements (Yifan et al., 2019) have incorporated surfels into a differentiable framework. However, these methods typically require ground truth (GT) geometry, depth sensor data, or operate under constrained scenarios with known lighting.
Inspired by these works, we propose 2D Gaussian Splatting for 3D scene reconstruction and novel view synthesis that combines the benefits of both worlds, while overcoming their limitations. Unlike 3DGS, our approach represents a 3D scene with 2D Gaussian primitives, each defining an oriented elliptical disk. The significant advantage of 2D Gaussian over its 3D counterpart lies in the accurate geometry representation during rendering. Specifically, 3DGS evaluates a Gaussian’s value at the intersection between a pixel ray and a 3D Gaussian (Keselman and Hebert, 2022, 2023), which leads to inconsistency depth when rendered from different viewpoints. In contrast, our method utilizes explicit ray-splat intersection, resulting in a perspective correct splatting, as illustrated in Figure 2, which in turn significantly improves reconstruction quality. Furthermore, the inherent surface normals in 2D Gaussian primitives enable direct surface regularization through normal constraints. In contrast with surfels-based models (Pfister et al., 2000; Zwicker et al., 2001a; Yifan et al., 2019), our 2D Gaussians can be recovered from unknown geometry with gradient-based optimization.
@@
  • We present a highly efficient differentiable 2D Gaussian renderer, enabling perspective-correct splatting by leveraging 2D surface modeling, ray-splat intersection, and volumetric integration.
@@ 2.1. Novel view synthesis
3D Reconstruction from multi-view images has been a long-standing goal in computer vision. Multi-view stereo based methods (Schönberger et al., 2016; Yao et al., 2018; Yu and Gao, 2020) rely on a modular pipeline that involves feature matching, depth prediction, and fusion. In contrast, recent neural approaches (Niemeyer et al., 2020; Yariv et al., 2020) represent surface implicitly via an MLP (Park et al., 2019; Mescheder et al., 2019) , extracting surfaces post-training via the Marching Cube algorithm. Further advancements (Oechsle et al., 2021; Wang et al., 2021; Yariv et al., 2021) integrated implicit surfaces with volume rendering, achieving detailed surface reconstructions from RGB images. These methods have been extended to large-scale reconstructions via additional regularization (Yu et al., 2022b; Li et al., 2023; Yu et al., 2022a), and efficient reconstruction for objects (Wang et al., 2023). Despite these impressive developments, efficient large-scale scene reconstruction remains a challenge. For instance, Neuralangelo (Li et al., 2023) requires 128 GPU hours for reconstructing a single scene from the Tanks and Temples Dataset (Knapitsch et al., 2017). In this work, we introduce 2D Gaussian splatting, a method that significantly accelerates the reconstruction process. It achieves similar or slightly better results compared to previous implicit neural surface representations, while being an order of magnitude faster.
@@ 2.4. Concurrent works
Challenges in Surface Reconstruction
Reconstructing surfaces using 3D Gaussian modeling and splatting faces several challenges. First, the volumetric radiance representation of 3D Gaussians conflicts with the thin nature of surfaces. Second, 3DGS does not natively model surface normals, essential for high-quality surface reconstruction. Third, the rasterization process in 3DGS lacks multi-view consistency, leading to varied 2D intersection planes for different viewpoints (Keselman and Hebert, 2023), as illustrated in Figure 2 (a). Additionally, using an affine matrix for transforming a 3D Gaussian into ray space only yields accurate projections near the center, compromising on perspective accuracy around surrounding regions (Zwicker et al., 2004). Therefore, it often results in noisy reconstructions, as shown in Figure 5.
@@ 4.1. Modeling
Unlike 3DGS (Kerbl et al., 2023), which models the entire angular radiance in a blob, we simplify the 3-dimensional modeling by adopting “flat” 2D Gaussians embedded in 3D space. With 2D Gaussian modeling, the primitive distributes densities within a planar disk, defining the normal as the direction of the steepest change of density. This feature enables better alignment with thin surfaces. While previous methods (Kopanas et al., 2021; Yifan et al., 2019) also utilize 2D Gaussians for geometry reconstruction, they require a dense point cloud or ground-truth normals as input. By contrast, we simultaneously reconstruct the appearance and geometry given only a sparse calibration point cloud and photometric supervision.
@@
(4) $P(u,v)=\mathbf{p}_{k}+s_{u}\mathbf{t}_{u}u+s_{v}\mathbf{t}_{v}v=\mathbf{H}(u,v,1,1)^{\mathrm{T}}$
@@ 4.2. Splatting
One common strategy for rendering 2D Gaussians is to project the 2D Gaussian primitives onto the image space using the affine approximation of the perspective projection (Zwicker et al., 2001b, a). However, as noted in  (Zwicker et al., 2004), this projection is only accurate at the center of the Gaussian and has increasing approximation error with increased distance to the center. To address this issue, Zwicker et al. proposed a formulation based on homogeneous coordinates. Specifically, projecting the 2D splat onto an image plane can be described by a general 2D-to-2D mapping in homogeneous coordinates. Let $\mathbf{W}\in{4\times 4}$ be the transformation matrix from world space to screen space. The screen space points are hence obtained by
(7) $\mathbf{x}=(xz,yz,z,1)^{\mathrm{T}}=\mathbf{W}P(u,v)=\mathbf{W}\mathbf{H}(u,v,1,1)^{\mathrm{T}}$
@@
Ray-splat Intersection
We efficiently locate the ray-splat intersections by finding the intersection of three non-parallel planes, a method initially designed for specialized hardware (Weyrich et al., 2007). Given an image coordinate $\mathbf{x}=(x,y)$, we parameterize the ray of a pixel as the intersection of two orthogonal planes: the x-plane and the y-plane. Specifically, the x-plane is defined by a normal vector $(-1,0,0)$ and an offset $x$. Therefore, the x-plane can be represented as a 4D homogeneous plane $\mathbf{h}_{x}=(-1,0,0,x)^{\mathrm{T}}$. Similarly, the y-plane is $\mathbf{h}_{y}=(0,-1,0,y)^{\mathrm{T}}$. Thus, the ray $\mathbf{x}=(x,y)$ is determined by the intersection of the $x$-plane and the $y$-planes.
Next, we transform both planes into the local coordinates of the 2D Gaussian primitives, the $uv$-coordinate system. Note that transforming points on a plane using a transformation matrix $\mathbf{M}$ is equivalent to transforming the homogeneous plane parameters using the inverse transpose $\mathbf{M}^{-\mathrm{T}}$. Therefore, applying $\mathbf{M}=(\mathbf{W}\mathbf{H})^{-1}$ is equivalent to $(\mathbf{W}\mathbf{H})^{\mathrm{T}}$, eliminating explicit matrix inversion and yielding:
(8) $\mathbf{h}_{u}=(\mathbf{W}\mathbf{H})^{\mathrm{T}}{\mathbf{h}_{x}}\quad\mathbf{h}_{v}=(\mathbf{W}\mathbf{H})^{\mathrm{T}}{\mathbf{h}_{y}}$
@@
(9) $\mathbf{h}_{u}\cdot(u,v,1,1)^{\mathrm{T}}=\mathbf{h}_{v}\cdot(u,v,1,1)^{\mathrm{T}}=0$
@@
where $\mathbf{h}_{u}^{i},\mathbf{h}_{v}^{i}$ are the $i$-th parameter of the 4D homogeneous plane parameters. We obtain the depth $z$ of the intersected points via Eq. 7 and evaluate the Gaussian value with Eq 6.
Degenerate Solutions
When a 2D Gaussian is observed from a slanted viewpoint, it degenerates to a line in screen space. Therefore, it might be missed during rasterization. To deal with these cases and stabilize optimization, we employ the object-space low-pass filter introduced in (Botsch et al., 2005):
@@
where $\mathbf{u}(\mathbf{x})$ is given by (10) and $\mathbf{c}$ is the projection of center $\mathbf{p}_{k}$. Intuitively, $\hat{\mathcal{G}}(\mathbf{x})$ is lower-bounded by a fixed screen-space Gaussian low-pass filter with center $\mathbf{c}$ and radius $\sigma$. In our experiments, we set $\sigma=\sqrt{2}/2$ to ensure sufficient pixels are used during rendering.
Rasterization
We follow a similar rasterization process as in 3DGS (Kerbl et al., 2023). First, a screen space bounding box is computed for each Gaussian primitive. Then, 2D Gaussians are sorted based on the depth of their center and organized into tiles based on their bounding boxes. Finally, volumetric alpha blending is used to integrate alpha-weighted appearance from front to back:
@@
Figure 4. Visual comparisons (test-set view) between our method, 3DGS (Kerbl et al., 2023), and SuGaR (Guédon and Lepetit, 2023) using scenes from an real-world dataset (Barron et al., 2022b). Our method excels at synthesizing geometrically accurate radiance fields and surface reconstruction, outperforming 3DGS and SuGaR in capturing sharp edges and intricate details.
@@ 5. Training
Depth Distortion
Different from NeRF, 3DGS’s volume rendering doesn’t consider the distance between intersected Gaussian primitives. Therefore, spreading out Gaussians might result in a similar color and depth rendering. This is different from surface rendering, where rays intersect the first visible surface exactly once. To mitigate this issue, we take inspiration from Mip-NeRF360 (Barron et al., 2022a) and propose a depth distortion loss to concentrate the weight distribution along the rays by minimizing the distance between the ray-splat intersections:
@@
2DGS-30k (Ours) 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80 10.9 m
Normal Consistency
As our representation is based on 2D Gaussian surface elements, we must ensure that all 2D splats are locally aligned with the actual surfaces. In the context of volume rendering where multiple semi-transparent surfels may exist along the ray, we consider the actual surface at the median point of intersection $\mathbf{p}_{s}$, where the accumulated opacity reaches 0.5. We then align the splats’ normal with the gradients of the depth maps as follows:
(14) $\mathcal{L}_{n}=\sum_{i}\omega_{i}(1-\mathbf{n}_{i}^{\mathrm{T}}\mathbf{N})$
where $i$ indexes over intersected splats along the ray, $\omega$ denotes the blending weight of the intersection point, $\mathbf{n}_{i}$ represents the normal of the splat that is oriented towards the camera, and $\mathbf{N}$ is the normal estimated by the gradient of the depth map. Specifically, $\mathbf{N}$ is computed with finite differences from nearby depth points as follows:
(15) $\mathbf{N}(x,y)=\frac{\nabla_{x}\mathbf{p}_{s}\times\nabla_{y}\mathbf{p}_{s}}{|\nabla_{x}\mathbf{p}_{s}\times\nabla_{y}\mathbf{p}_{s}|}$
@@
Final Loss
Finally, we optimize our model from an initial sparse point cloud using a set of posed images. We minimize the following loss function:
@@
where $\mathcal{L}c$ is an RGB reconstruction loss combining $\mathcal{L}_{1}$ with the D-SSIM term from (Kerbl et al., 2023), while $\mathcal{L}_{d}$ and $\mathcal{L}_{n}$ are regularization terms. We set $\alpha=1000$ for bounded scenes, $\alpha=100$ for unbounded scenes, and $\beta=0.05$ for all scenes.